# Spiking GS: Towards High-Accuracy and Low-Cost Surface Reconstruction via Spiking Neuron-based Gaussian Splatting

### v3

@@ 4.1 FIF Neurons in Spiking GS
Global FIF neuron on opacity. The global FIF neuron is applied to the Gaussians’ opacity $\alpha$. The output of FIF neuron is calculated as:
@@
where $\bar{V}^{\alpha}$ is shared by all Gaussians and optimized during training. In the original method, a key criterion for pruning Gaussians is whether their opacity is below a predefined threshold $\epsilon_{\alpha}$. Additionally, a reset strategy is employed, setting all Gaussians’ opacities to a low value $\epsilon^{\mathrm{r}}_{\alpha}$, together with pruning to manage the increasing number of Gaussians. Although this approach is effective in reducing floaters near the camera, it is less successful in reducing the number of LOGs around surfaces. By replacing $\epsilon_{\alpha}$ and $\epsilon^{\mathrm{r}}_{\alpha}$ with $\bar{V}^{\alpha}$, Gaussians with $\alpha<\bar{V}^{\alpha}$, typically corresponding to LOGs, are efficiently removed. In addition, the use of a larger adaptive reset threshold can implicitly guide 3DGS to converge toward the optimal case across various scenes.
Local FIF neuron on representation function. The tails of Gaussians (low function value parts) often overlap when they are positioned closely together. This overlap can sometimes contribute to the integration for view rendering and surface reconstruction, resulting in biases and an excessive number of Gaussians. To address this, we “cut off” the tails of the 3D Gaussian representation function to reduce such overlap. This is accomplished by introducing the local FIF (full-precision integrate-and-fire) neuron into the Gaussian representation function, redefined as:
@@
where $\bar{V}^{\mathrm{p}}$ is the cut-off threshold set as a learnable parameter. Notably, each Gaussian is assigned a unique $\bar{V}^{\mathrm{p}}_{i}$, reflecting the fact that different regions in the scene may need varying cut-off thresholds to capture the geometry accurately. For instance, areas near object boundaries might necessitate a larger cut-off threshold to model the discontinuity at the boundary than other parts of the object. We find the local FIF neurons effectively reduce the contribution of low-opacity tails (LOTs), minimizing unwanted overlap and improving the accuracy and efficiency of the 3DGS. In practice, to better integrate $\hat{\mathcal{G}}$ into the rasterization framework, we directly apply the local FIF neuron on the projected 2D Gaussian on the screen, making it act as if applying the model on the 3D Gaussian.
@@ 4.2 Optimizing Spiking GS
where $\lambda_{\alpha}=\lambda_{\mathrm{p}}=2e^{-5}$.
Scale loss. As presented in Fan et al. (2024), large-sized Gaussians tend to magnify minor errors during training and hinder the Gaussians’ ability to capture high-frequency geometric details. Motivated by this insight, we propose a simple scale loss to constrain the scale of Gaussians:
@@
Total loss. Excluding the losses on $\bar{V}^{\alpha}$, $\bar{V}^{\mathrm{p}}$, and $\mathbf{S}$, we implement losses of previous methods. Except for the photometric loss $\mathcal{L}_{c}$ following Kerbl et al. (2023), we also use depth distortion loss $\mathcal{L}_{d}$, normal loss $\mathcal{L}_{N}$ following Huang et al. (2024), and total variance loss $\mathcal{L}_{t}$ following Karnieli et al. (2022) to regularize the geometry. To sum up, the total loss $\mathcal{L}$ is calculated as:
@@
Scale-based Clone. The geometry of blind spots or regions observed by limited viewpoints is particularly difficult to reconstruct due to the lack of constraints from viewpoints. Additionally, we find Gaussians in those regions hardly meet the clone criteria of splitting and cloning, which may be overly removed by the enhanced pruning strategy, resulting in deficiency (e.g., undesired holes and pits) in the reconstructed surface. To mitigate this issue, we propose a scale-based clone. Specifically, it additionally clone Gaussians that meet $\mathcal{R}\left(\text{max}(\mathbf{S}^{\mathrm{x}}_{i},\mathbf{S}^{\mathrm{y}}_{i})\right)\in[\theta-\delta,\theta+\delta]$, where $\delta=\frac{\theta}{200}$. The new clone rule enriches the number of Gaussians in blind spot regions, improves the opportunity for those Gaussians to be guided by view-related loss, compensates for the geometric information deficiency, and improves the reconstruction quality in these regions.
@@ 5 Experiments
Our training framework is developed on 3DGS Kerbl et al. (2023) with flattened Gaussians Dai et al. (2024) as primitive and customized CUDA kernels [^1] For parameters setup and training details, please refer to the appendix..
Evaluation metrics and datasets. To evaluate the accuracy of the 3D surface reconstruction, we use the Chamfer distance as the metric Dai et al. (2024); Huang et al. (2024); Liao et al. (2024). We also report training time and the number of Gaussians (noted as #G) to evaluate the efficiency of our method. We test our method in 8 scenes from NeRF-Synthetic dataset Mildenhall et al. (2021), 6 scenes from Dex-NeRF dataset Ichnowski et al. (2021), and 15 scenes from DTU dataset Jensen et al. (2014). These datasets include both synthetic and real-world scenes.
Baselines. We compare our method’s performance with one NeRF-based neural surface reconstruction method (Spiking NeRF Liao et al. (2024), noted as “SpNeRF”) and five Gaussian-based surface reconstruction methods (i.e, 3DGS Kerbl et al. (2023), SuGaR Guédon & Lepetit (2024), G-Surfel Dai et al. (2024), 2DGS Huang et al. (2024), and GOF Yu et al. (2024)). To further evaluate the training and storage efficiency, we compare our method with two compression methods (i.e, Compact-3DGS Lee et al. (2024), denoted as “C3DGS”, and HAC Chen et al. (2024c)).
Figure 4: Qualitative Comparisons of surface reconstruction performed on DTU Jensen et al. (2014) and Dex-NeRF datasets Ichnowski et al. (2021). We show the Chamfer distance in the bottom left corner of the image.
@@ 5.1 Geometry Reconstruction
Quantitative Comparison. To quantitatively assess the quality of the reconstructed surface, we compared the proposed Spiking GS with other methods on the NeRF-Synthetic Mildenhall et al. (2021), Dex-NeRF Ichnowski et al. (2021), and DTU Jensen et al. (2014) datasets. The results are shown in Tab. 1-3. respectively. According to the results, our method achieves SOTA performance on most datasets (NeRF-Synthetic and Dex-NeRF). We find that the adaptability of the FIF neurons improves the method’s robustness, showing stability in various scenes, especially with semi-transparent objects. However, we also notice that the advantage diminishes on the DTU dataset primarily because the viewpoints in the DTU are concentrated on the front side of the object, with few views from the back. This results in limited constraints introduced from the back views, negatively impacting our method.
@@
Table 1: Quantitative Comparison on NeRF-Synthetic dataset Mildenhall et al. (2021). We report the Chamfer distance (× $\times\ 10^{-2}$) for the reconstructed mesh in 8 scenes, the number of Gaussians (#G), and training time.
@@
- w/o glb 0.46 1.48 0.74 1.12 0.88 0.98 0.64 1.02 0.92 238k 12.6 m
- w/o loc 0.47 1.37 0.70 1.10 0.81 0.95 0.60 1.00 0.88 85k 10.3 m
Table 2: Quantitative Comparison on selected scenes of Dex-NeRF dataset Ichnowski et al. (2021). We report the Chamfer distance (× $\times\ 10^{-2}$), the number of Gaussians (#G), and training time.
@@
Table 3: Quantitative Comparison on $15$ scenes from DTU dataset Jensen et al. (2014). We report the Chamfer distance, the number of Gaussians (#G), and training time.
  Method 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean #G Time
  SpNeRF 0.84 1.20 1.02 0.38 1.15 0.72 0.69 1.10 1.19 0.65 0.49 1.60 0.49 0.55 0.51 0.83 - $>$  12h
3DGS 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96 542k 11.2m
SuGaR 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33 - ~ 1h
G-Surfel 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88 166k 16.5m
2DGS 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80 224k 18.8m
GOF 0.52 0.91 0.40 0.37 1.11 0.88 0.72 1.18 1.26 0.75 0.72 0.88 0.46 0.71 0.55 0.76 369k 1.8h
  Ours 0.63 0.80 0.50 0.46 1.04 1.01 0.87 1.15 1.32 0.61 0.73 1.19 0.44 0.68 0.48 0.79 171k 11.1m
- w/o glb 0.69 0.86 0.68 0.48 1.17 1.27 0.92 1.19 1.35 0.63 0.76 1.21 0.50 0.76 0.54 0.87 427k 15.9m
- w/o loc 0.73 0.81 0.68 0.49 1.04 1.02 0.93 1.22 1.31 0.60 0.77 1.28 0.45 0.69 0.49 0.83 191k 11.2m
@@
Table 4: Performance comparion between C3DGS Lee et al. (2024), HAC Chen et al. (2024c), and Spiking GS (ours) on three datasets Ichnowski et al. (2021); Jensen et al. (2014); Mildenhall et al. (2021). We report the averaged Chamfer distance (CD), the number of Gaussians (#G), and reconstruction time. Bold numbers indicate the best results.
@@
Figure 5: The qualitative comparison between the Spiking GS and its alternatives. From left to right: columns 1-3 are comparisons of reconstructed mesh between the full model and ‘w/o glb’, and columns 4-6 are comparisons of Gaussians’ center points projected on a specific view between the full model and ‘w/o loc’.
Efficiency Comparison. As shown in Tables 1-3, our method demonstrates a significant reduction in storage (nearly twice fewer Gaussians than 2DGS Yu et al. (2024) on the NeRF-Synthetic and DTU dataset, and even five times fewer than 2DGS Yu et al. (2024) on the Dex-NeRF dataset) while maintaining a competitively fast training time. Additionally, we also compare Spiking GS with other compression methods (C3DGS Lee et al. (2024) and HAC Chen et al. (2024c)) in terms of Chamfer distance, number of Gaussians, and training time. The results are shown in Tab. 4, indicating our method balances the reconstruction accuracy, training efficiency, and storage efficiency. In practice, many factors may affect the method’s efficiency. For instance, although reducing the number of Gaussians during training can shorten training time, the additional operations (e.g., regularization computation) required to achieve this may increase time, offsetting the benefits. This explains why our method is faster than C3DGS and HAC, despite sometimes using more Gaussians. Additionally, the number of training epochs plays a critical role. For example, with a pre-trained normal estimation model, G-Surfel requires only 15000 iterations for good performance. In contrast, our method and other 3DGS approaches, lacking such priors, need more iterations.
5.2 Ablation Study
To validate the effectiveness of the proposed methods, we conduct an ablation study of the proposed Spiking GS. The results of a comparison between our method and two alternatives, ‘w/o glb’ and ‘w/o loc’, are shown in Tab.1  [^2] More results can be found in the appendix.. Specifically, ‘w/o glb’ removes the global FIF neuron, while ‘w/o loc’ excludes the local FIF neuron. The results show that omitting either the global or local FIF neurons negatively impacts both the accuracy and efficiency of surface reconstruction. The two types of FIF neurons play distinct roles in our method for reducing LOPs, resulting in different effects. For instance, the global FIF neuron is primarily effective in reducing the number of LOGs to improve reconstruction quality and efficiency. Particularly, we find method with global FIF neuron can better capture the details on the surface, validated by a quality comparison between the full model and ‘w/o glb’ in Fig. 5, column 1-3. In contrast, local FIF neurons decrease LOTs to minimize the overlap of Gaussians, resulting in more separated centers of Gaussians, as evidenced by the visualized Gaussian center’s distribution in Fig. 5, columns 5-6. Local FIF neurons are particularly effective in specific scenes, such as scenes with flat surfaces (the round table in Fig. 4, column 2) where tail overlap is predominant. Incorporating local FIF neurons will be more efficient in these cases to reduce the number of Gaussians, compared to using global FIF neurons.

### v5

@@ 4.1 FIF Neurons in Spiking GS
Global FIF neuron on opacity. The global FIF neuron is implemented on the Gaussians’ opacity $\alpha$. The output of FIF neuron is calculated as:
@@
where $\bar{V}^{\alpha}$ is shared by all Gaussians and optimized during training. Previous 3DGS methods [18, 14] control the density of Gaussians by pruning and opacity resetting. Specifically, Gaussians with opacity values below a threshold $\epsilon_{\alpha}$ are removed during optimization, and their opacities are periodically reset to a low value, $\epsilon^{\mathrm{r}}_{\alpha}$. Although this density control can reduce LOGs near the camera, it is less effective at managing LOGs around surfaces. In contrast, the proposed global FIF neuron leverages the discontinuity and differentiability of spiking neurons equipped with surrogate gradient to provide a more adaptive control over LOGs during training. Specifically, Gaussians are set to zero opacity when their values fall below a learnable threshold $\bar{V}^{\alpha}$. These Gaussians, often corresponding to LOGs, are then removed in pruning. Additionally, we replace $\epsilon^{\mathrm{r}}_{\alpha}$ with $\bar{V}^{\alpha}$ in opacity-resetting, where $\bar{V}^{\alpha}$ is consistently greater than $\epsilon^{\mathrm{r}}_{\alpha}$. This higher opacity-resetting threshold increases the overall opacity in scenes, which implicitly guides 3DGS to generate less LOGs.
Local FIF neuron on representation function. The tails of Gaussians (low function value parts, i.e, LOTs) often overlap when they are positioned closely together. This overlap can sometimes contribute to the integration for view rendering and surface reconstruction, resulting in biases and an excessive number of Gaussians. To eliminate the contribution of LOTs, we use the local FIF neuron to set the value of the low-function-value parts to zero. Specifically, the Gaussian’s representation function is defined as:
@@
where $\bar{V}^{\mathrm{p}}$ is the cut-off threshold set as a learnable parameter. Notably, each Gaussian is assigned a unique, learnable $\bar{V}^{\mathrm{p}}_{i}$, reflecting the fact that different regions in the scene may need varying cut-off thresholds to capture the geometry accurately. For instance, areas near object boundaries might necessitate a larger cut-off threshold to model the discontinuity at the boundary than other parts of the object. In practice, to better integrate $\hat{\mathcal{G}}$ into the rasterization framework, we directly apply the local FIF neuron on the projected 2D Gaussian on the screen, making it act as if applying the model on the 3D Gaussian.
@@ 4.2 Optimizing Spiking GS
where $\lambda_{\alpha}=\lambda_{\mathrm{p}}=2e^{-5}$. We experimentally find this hyperparameter setup fit general cases. Large values will cause Gaussians to be overly removed. Overly small value weakens the influence of spiking neurons.
Scale loss. As presented in [9], large-sized Gaussians tend to magnify minor errors during training and hinder the Gaussians’ ability to capture high-frequency geometric details. Motivated by this insight, we propose a simple scale loss to constrain the scale of Gaussians:
@@
Total loss. Excluding the losses on $\bar{V}^{\alpha}$, $\bar{V}^{\mathrm{p}}$, and $\mathbf{S}$, we implement losses of previous methods. Except for the photometric loss $\mathcal{L}_{c}$ following [18], we also use depth distortion loss $\mathcal{L}_{d}$, normal loss $\mathcal{L}_{N}$ following [14], and total variance loss $\mathcal{L}_{t}$ following [17] to regularize the geometry. To sum up, the total loss $\mathcal{L}$ is calculated as:
@@
Scale-based Clone. The geometry of blind spots or regions observed by limited viewpoints is particularly difficult to reconstruct due to the lack of constraints from viewpoints. Additionally, we find Gaussians in those regions hardly meet the clone criteria of splitting and cloning, which may be overly removed by the enhanced pruning strategy, resulting in deficiency (e.g., undesired holes and pits) in the reconstructed surface. To mitigate this issue, we propose a scale-based clone. Specifically, it additionally clone Gaussians that meet $\text{max}(\mathbf{S}^{\mathrm{x}}_{i},\mathbf{S}^{\mathrm{y}}_{i})\in[\bar{V}^{\theta}-\delta,\bar{V}^{\theta}+\delta]$, where $\delta=\frac{\bar{V}^{\theta}}{200}$. The rationale behind this approach is that blind-spot regions lack view constraints, resulting in small gradients in view-space positions. Consequently, Gaussians in these regions are often not cloned appropriately. We select $\bar{V}^{\theta}$ (the scale loss’s hyperparameter) since the scale loss is view-invariant and becomes the dominant term in these regions. Thus, Gaussians with scales near this value are likely to be located in blind-spot regions. The new clone rule enriches the number of Gaussians in these regions, and mitigates artifacts in reconstructed such as holes.
@@ 5 Experiments
Figure 6: Qualitative Comparisons of surface reconstruction performed on DTU [16] and Dex-NeRF datasets [15]. We show the Chamfer distance in the bottom left corner of the image.
Our training framework is developed on 3DGS [18] with flattened Gaussians [5] as primitive [^2] For parameters setup and training details, please refer to the appendix..
Evaluation metrics and datasets. We test our method in 8 scenes from NeRF-Synthetic dataset [29], 6 scenes from Dex-NeRF dataset [15], and 15 scenes from DTU dataset [16]. These datasets include both synthetic and real-world scenes. To evaluate the accuracy of the 3D surface reconstruction, we use the Chamfer distance as the metric [5, 14, 25]. To evaluate the rendering quality, we use the standard PSNR, SSIM, and LPIPS metrics and report the results on NeRF-Synthetic and Dex-NeRF datasets on the test set. We also report training time and Gaussians’ number (noted as #G) to evaluate the efficiency of our method.
Baselines. We compare our method’s performance with one NeRF-based neural surface reconstruction method (Spiking NeRF [25], noted as “SpNeRF”) and five Gaussian-based surface reconstruction methods (i.e, 3DGS [18], SuGaR [12], G-Surfel [5], 2DGS [14], and GOF [41]). To further evaluate the training and storage efficiency, we compare our method with two compression methods (i.e, Compact-3DGS [21], denoted as “C3DGS”, and HAC [4]).
@@ 5.1 Geometry Reconstruction
Quantitative Comparison. To quantitatively assess the quality of the reconstructed surface, we compared the proposed Spiking GS with other methods on the NeRF-Synthetic [29], Dex-NeRF [15], and DTU [16] datasets. The results are shown in Tab. 1-3. respectively. According to the results, our method achieves SOTA performance on most datasets (NeRF-Synthetic and Dex-NeRF). We find that the adaptability of the FIF neurons improves the method’s robustness, showing stability in various scenes, especially with semi-transparent objects. However, we also notice that the advantage diminishes on the DTU dataset primarily because the viewpoints in the DTU are concentrated on the front side of the object, with few views from the back. This results in limited constraints introduced from the back views, negatively impacting our method.
@@
Table 1: Quantitative Comparison on NeRF-Synthetic dataset [29]. We report the Chamfer distance (× $\times\ 10^{-2}$) for the reconstructed mesh in 8 scenes, the number of Gaussians (#G), and training time. Best results are highlighted as 1st , 2nd , and 3rd .
@@
Table 2: Quantitative Comparison on selected scenes of Dex-NeRF dataset [15]. We report the Chamfer distance (× $\times\ 10^{-2}$), the number of Gaussians (#G), and training time.
@@
Table 3: Quantitative Comparison on $15$ scenes from DTU dataset [16]. We report the Chamfer distance, the number of Gaussians (#G), and training time.
  SpNeRF 3DGS SuGaR G-Surfel 2DGS GOF Ours
  24 0.84 2.14 1.47 0.66 0.48 0.52 0.63
37 1.20 1.53 1.33 0.93 0.91 0.91 0.80
40 1.02 2.08 1.13 0.54 0.39 0.40 0.50
55 0.38 1.68 0.61 0.41 0.39 0.37 0.46
63 1.15 3.49 2.25 1.06 1.01 1.11 1.04
65 0.72 2.21 1.71 1.14 0.83 0.88 1.01
69 0.69 1.43 1.15 0.85 0.81 0.72 0.87
83 1.10 2.07 1.63 1.29 1.36 1.18 1.15
97 1.19 2.22 1.62 1.53 1.27 1.26 1.32
105 0.65 1.75 1.07 0.79 0.76 0.75 0.61
106 0.49 1.79 0.79 0.82 0.70 0.72 0.73
110 1.60 2.55 2.45 1.58 1.40 0.88 1.19
114 0.49 1.53 0.98 0.45 0.40 0.46 0.44
118 0.55 1.52 0.88 0.66 0.76 0.71 0.68
122 0.51 1.50 0.79 0.53 0.52 0.55 0.48
  AVG 0.83 1.96 1.33 0.88 0.80 0.76 0.79
#G - 542k - 166k 224k 369k 171k
Time $>$  12h 11.2m ~ 1h 16.5m 18.8m 1.8h 11.1m
Figure 7: The qualitative comparison between the Spiking GS and its alternatives. Columns 1-3 are the comparison of reconstructed mesh between the full model and ‘w/o glb’. Columns 4-6 are the comparison of Gaussians’ center points projected on a specific view between the full model and ‘w/o loc’. Columns 7-9 are the mesh comparison between the full model and ‘w/o scale-based clone’.
Efficiency Comparison. As shown in Tables 1-3, our method demonstrates a significant reduction in storage (nearly twice fewer Gaussians than 2DGS [41] on the NeRF-Synthetic and DTU dataset, and even five times fewer than 2DGS [41] on the Dex-NeRF dataset) while maintaining a competitively fast training time. Additionally, we also compare Spiking GS with other compression methods (C3DGS [21] and HAC [4]) in terms of Chamfer distance, number of Gaussians, and training time. The results are shown in Tab. 5, indicating our method balances the reconstruction accuracy, training efficiency, and storage efficiency. In practice, many factors may affect the method’s efficiency. For instance, although reducing the number of Gaussians during training can shorten training time, the additional operations (e.g., regularization computation) required to achieve this may increase time, offsetting the benefits. This explains why our method is faster than C3DGS and HAC, despite sometimes using more Gaussians. Additionally, the number of training epochs plays a critical role. For example, with a pre-trained normal estimation model, G-Surfel requires only 15000 iterations for good performance. In contrast, our method and other 3DGS approaches, lacking such priors, need more iterations.
5.2 Novel View Rendering
We present the results of the novel view rendering in Tab. 4. Generally considering three evaluation metrics, our method shows comparable performance to G-Surfel [5], while perform slightly worse than GOF [41]. We believe such a downgrading in rendered views’ quality is acceptable as we use much less Gaussians to represent the scene. It also worth mentioning our method reach a superior performance in geometry reconstruction.
5.3 Ablation Study
In this section, we show quantitative analyses regarding the important contributions of Spiking GS (i.e, the effect of global and local FIF neurons and the scale-based clone) and used regularization losses. This results in a comparison between our method and six alternatives, as shown in Tab.6.
The results show that omitting either the global (‘w/o glb’) or local FIF (‘w/o loc’) neurons negatively impacts both the accuracy and efficiency of surface reconstruction. The two types of FIF neurons play distinct roles in our method for reducing LOPs, resulting in different effects. For instance, the global FIF neuron is primarily effective in reducing the number of LOGs to improve reconstruction quality and efficiency. Particularly, we find a method with the global FIF neuron can better capture the details on the surface, validated by a quality comparison between the full model and ‘w/o glb’ in Fig. 7, columns 1-3. In contrast, local FIF neurons decrease LOTs to minimize the overlap of Gaussians, resulting in more separated centers of Gaussians, as evidenced by the visualized Gaussian center’s distribution in Fig. 7, columns 4-6. Local FIF neurons are particularly effective in specific scenes, such as scenes with flat surfaces (the round table in Fig. 6, row 2) where tail overlap is predominant. Incorporating local FIF neurons will be more efficient in these cases to reduce the number of Gaussians, compared to using global FIF neurons. Removing scale-based clone (‘w/o c’) will cause holes-like artifacts in reconstructed surface, as shown in Fig. 7, column 9, which harms the accuracy of surface reconstruction. Moreover, the normal (‘w/o $\mathcal{L}_{N}$’), depth distortion (‘w/o $\mathcal{L}_{d}$’), and total variance loss (‘w/o $\mathcal{L}_{t}$’) are important empirical priors to enhance surface reconstruction quality [14, 5], which will harm the performance if we remove them. We would like to emphasize that including these priors does not diminish our contributions. While previous methods focus on refining priors to improve accuracy, they overlook the inherent deficiencies in 3DGS representations (i.e, excessive LOPs). The proposed Spiking GS improve the accuracy and efficiency of surface reconstruction in a different, novel perspective, by identifying and addressing the LOPs through spiking neurons.
Table 4: PSNR, SSIM, and LPIPS scores for NeRF-Synthetic [29] and Dex-NeRF [15] datasets.
  Dataset NeRF-Synthetic Dex-NeRF
[…]
  3DGS 31.05 0.959 0.051 21.84 0.904 0.287
G-Surfel 31.10 0.954 0.050 30.77 0.953 0.123
2DGS 30.28 0.956 0.054 22.19 0.886 0.280
GOF 33.78 0.969 0.031 20.01 0.913 0.241
  Ours 30.23 0.954 0.050 30.59 0.953 0.111
@@ 5.3 Ablation Study
Table 5: Performance comparion between C3DGS [21], HAC [4], and Spiking GS (ours) on three datasets [15, 16, 29]. We report the averaged Chamfer distance (CD), the number of Gaussians (#G), and reconstruction time. Bold numbers indicate the best results.
@@
Table 6: Ablation study on NeRF-Synthetic dataset [29]. We compare the proposed Spiking GS (Full) with its 6 alternatives (i.e, ‘w/o glb’, ‘w/o loc’, ‘w/o c’, ‘w/o $\mathcal{L}_{N}$’, ‘w/o $\mathcal{L}_{d}$’, ‘w/o $\mathcal{L}_{t}$’).
  Full w/o glb w/o loc w/o c w/o $\mathcal{L}_{N}$ w/o $\mathcal{L}_{d}$ w/o $\mathcal{L}_{t}$
[…]